# Empower Structure-Based Molecule Optimization with Gradient Guidance

### v3

@@ Abstract
Structure-based molecule optimization (SBMO) aims to optimize molecules with both continuous coordinates and discrete types against protein targets. A promising direction is to exert gradient guidance on generative models given its remarkable success in images, but it is challenging to guide discrete data and risks inconsistencies between modalities. To this end, we leverage a continuous and differentiable space derived through Bayesian inference, presenting Molecule Joint Optimization (MolJO), the gradient-based SBMO framework that facilitates joint guidance signals across different modalities while preserving SE(3)-equivariance. We introduce a novel backward correction strategy that optimizes within a sliding window of the past histories, allowing for a seamless trade-off between explore-and-exploit during optimization. MolJO achieves state-of-the-art performance on CrossDocked2020 benchmark (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), more than 4×4$\times$4 × improvement in Success Rate compared to the gradient-based counterpart, and 2×2$\times$2 × “Me-Better” Ratio as much as 3D baselines. Furthermore, we extend MolJO to a wide range of optimization settings, including multi-objective optimization and challenging tasks in drug design such as R-group optimization and scaffold hopping, further underscoring its versatility.
@@ 1 Introduction
Structure-based drug design (SBDD) plays a critical role in drug discovery by identifying three-dimensional (3D) molecules that are favorable to protein targets (Isert et al., 2023). While recent SBDD focuses on the initial identification of potential drug candidates, these compounds must undergo a series of further modifications for optimized properties, a process that is both complex and time-consuming (Hughes et al., 2011). Therefore, structure-based molecule optimization (SBMO) has garnered increasing interest in real-world drug design (Zhou et al., 2024a), emphasizing the practical need to optimize 3D molecules to meet specific therapeutic criteria.
@@
In this paper, we address the multi-modality challenge for gradient guidance by leveraging a continuous and differential space, representing an aggregation of noisy samples from the data space derived through Bayesian inference (Graves et al., 2023). We design MolJO (Molecule Joint Optimization), a principled, ent-to-end differentiable framework that enables gradient-based optimization of continuous and discrete variables. We introduce a novel sampling strategy called backward correction, enhancing the alignment of gradients over different steps. By maintaining a sliding window of past history for optimization, the backward correction strategy enforces explicit dependency on the past, effectively alleviating the issue of inconsistencies. Moreover, it balances the exploration of molecular space with the exploitation of better-aligned guidance signals, offering a flexible trade-off.
@@ 6 Conclusion
We present MolJO, the joint gradient-based SE(3)-equivariant framework to solve the structure-based molecule optimization problems, which only requires differentiable energy functions instead of expensive oracle simulations. The general framework further equips gradient-based optimization method with backward correction strategy, offering a flexible trade-off between exploration and exploitation. Experiments show that MolJO is able to improve the binding affinity of molecules by establishing more key interactions and enhance drug-likeness and synthesizability, achieving state-of-the-art performance on CrossDocked2020 (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), together with 4×4$\times$4 × improvement compared to gradient-based counterpart and 2×2$\times$2 × “Me-Better” Ratio as other 3D baselines.
@@ Acknowledgments
This work is supported the Natural Science Foundation of China (Grant No. 62376133) and sponsored by Beijing Nova Program (20240484682) and the Wuxi Research Institute of Applied Technologies, Tsinghua University (20242001120).
@@ Appendix A Overview of Bayesian Flow Networks
In this section, we provide a further explanation of Bayesian Flow Networks (BFNs) that are designed to model the generation of data through a process of message exchange between a “sender” and a “receiver” (Graves et al., 2023). The fundamental elements include the sender distribution for the sender, and input distribution, output distribution, receiver distribution for the receiver.
@@
This process is framed within the context of Bayesian inference, where the sender distribution is a factorized distribution $p_{S}(\mathbf{y}|\mathbf{m},\alpha_{t})$ that introduces noise to each dimension of the data $\mathbf{m}$ and sends it to the receiver, which observes the $\mathbf{y}$ and has access to the noisy channel with accuracy $\alpha$ at timestep $t$, and compares it with its own receiver distribution $p_{R}(\mathbf{y}|\bm{\theta},\mathbf{p};t)$ based on its current belief of the parameters $\bm{\theta}$, the timestep and any conditional input such as protein pocket $\mathbf{p}$.
@@
The generative process for the receiver begins with a prior distribution (referred to as input distribution $p_{I}(\mathbf{m}|\bm{\theta})=\prod_{d=1}^{N}p_{I}(m^{(d)}|\theta^{(d)})$ that is also factorized for $N$-dimensional data) that defines its initial belief about the data. For continuous data, the prior can be chosen as a Gaussian (Song et al., 2024; Qu et al., 2024), or other distribution such as von Mises distribution (Wu et al., 2025), while for discrete data, the prior is modeled as a uniform categorical distribution. Then, the receiver uses its belief $\bm{\theta}$ with the help of neural network to model the inter-dependency among dimensions and compute the output distribution $p_{O}(\hat{\mathbb{m}}|\bm{\theta},\mathbf{p};t)$, which represents its estimate of the possible reconstructions of original data $\mathbf{m}$, and is used to construct the receiver distribution (Eq. 2).
@@
The Bayesian update function in Eq. 1 defines how the prior belief $\bm{\theta}_{0}$ is updated to the conjugate posterior $\bm{\theta}_{t}$. Ideally, the update requires aggregating all possible noisy latents $\mathbf{y}$ from the sender distribution $p_{S}$, while during the actual generative process, there is only receiver distribution $p_{R}$, leaving an exposure bias, and different approximations determine different forms of mapping $\theta=f(\mathbf{y})$ to the posterior, showcasing the flexibility in the design space of BFN.
@@
Through the iterative communication between sender and receiver, the receiver progressively updates its belief of the underlying parameters, and training is achieved by minimizing the divergence between the sender and receiver distributions (Eq. 3). This is analogous to the way Bayesian inference works in parameter estimation: as more noisy data $\mathbf{y}$ is observed, the receiver’s posterior belief about the data $\mathbf{m}$ becomes increasingly accurate, which implies the reconstruction would be made easier. It is shown that BFN for Gaussian priors can be sampled from the view of SDE (Xue et al., 2024), but this kind of sampling displays distinct behaviors empirically, possibly due to discretization errors.
@@ Appendix B Sender Distribution for Discrete Data
While true discrete data can be viewed as a sharp one-hot distribution, it can be further relaxed by a factor $\omega\in[0,1]$ into a Categorical distribution, with the probability $p(k^{(d)}|\mathbf{v}^{(d)};\omega)=\frac{1-\omega}{K}+\omega\delta_{k^{(d)}\mathbf{v}^{(d)}}$ for $k$ from $1$ to $K$ along the $d$-th dimension, where $\delta$ is the Kronecker delta function.
Instead of focusing on the density or sampling once from it, note that the counts $c$ of observing each class in $m$ independent draws follow a multinomial distribution, namely $c\sim\text{Multi}(m,p)$. Dropping the superscripts, Graves et al. (2023) derives the following conclusions:
@@
Thus it naturally follows that such noisy $\mathbf{y}^{v}\sim\mathcal{N}(\mathbf{y}^{v}|\alpha(K\mathbf{e}_{\mathbf{v}}-\mathbb{1}),\alpha K\mathbb{I})$.
@@ C.1 Proof of Guided Bayesian Update Distribution
where we have omitted some parentheses and protein pocket condition $\mathbf{p}$ for brevity.
@@
Eq. 19 together with the lemma above leads to Proposition 4.1.
@@ C.2 Proof of Equivariance
For any orthogonal matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ such that $\mathbf{R}^{\top}\mathbf{R}=\mathbb{I}$, it is easy to see that the prior $\bm{\theta}^{x}_{0}=\mathbb{0}$ is O(3)-invariant. Given that $\hat{\mathbf{x}}\sim p_{O}(\hat{\mathbf{x}}~{}|~{}\bm{\Phi}(\bm{\theta},\mathbf{p},t))$ and the equivariance of $\bm{\Phi}$, it suffices to prove the invariant likelihood for the transition kernel to complete the proof.
Given the parameterization of pretrained energy function $E(\bm{\theta},\mathbf{p},t)$ is SE(3)-equivariant, then the gradient $\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta})=-\nabla_{\bm{\theta}^{x}}E(\bm{\theta},\mathbf{p},t_{i})$ is also equivariant according to Bao et al. (2022).
@@
It also applies to cases where $i>k$, for we can recurrently view the starting point of backward corrected history $\bm{\theta}^{x}_{i-k}$ as the new O(3)-invariant prior $\bm{\theta}^{x}_{0}$ and iteratively make the above derivation. ∎
@@ D.1 Model Details
In order for a differentiable oracle function, we additionally train the energy function based on the molecules and their properties (Vina Score, QED, SA) in CrossDocked dataset (Francoeur et al., 2020) by minimizing the squared loss for property $c$ over the data distribution $p_{\mathrm{data}}$:
@@
The input parameters to all energy functions belong to the parameter space defined by $\beta_{1}=1.5$ for atom types, $\sigma_{1}=0.03$ for atom coordinates, $n=1000$ discrete steps. The energy network is parameterized with the same model architecture as TargetDiff (Guan et al., 2022), i.e. kNN graphs with $k=32$, $N=9$ layers with $d=128$ hidden dimension, 16-headed attention, and the same featurization, i.e. protein atoms (H, C, N, O, S, Se) and ligand atoms (C, N, O, F, P, S, Cl). For training, Adam optimizer is adopted with learning rate 0.0005, batch size is set to 8. The training takes less than 8 hours on a single RTX 3090 and converges within 5 epochs.
@@
To sample via guided Bayesian flow, we set the sample steps to 200, and the guidance scale to 50. For the combination different objectives, we merely take an average of different gradients.
@@ D.2 Task Details
Cases of lead optimization involve retaining the scaffold while redesigning the remaining R-groups, usually when the scaffold forms desirable interactions with the protein and anchors the binding mode, and the remaining parts need further modifications to secure this pattern and enhance binding affinity. Following Polykovskiy et al. (2020), we employ RDKit for fragmentization and atom annotation with R-group or Bemis-Murcko scaffold.
@@
Different from R-group design, scaffold hopping means to redesign the scaffold for a given molecule while keeping its core functional groups, for example to overcome the patent protection for a known drug molecule while retaining pharmaceutical activity. This is a technically more challenging task for generative models, for the missing parts they need to fulfill are generally larger than those in R-group design, and the hopping is usually subject to more chemical constraints. We construct scaffold hopping as a dual problem to R-group optimization with Bemis-Murcko scaffolding annotation, although it does not need to be so.
@@ Appendix E Effect of Molecular Size on Properties
The size of molecules are found with a notable impact on molecular properties, including Vina affinities (Qu et al., 2024). We quantify the relationship and plot the distribution of molecular properties w.r.t. the number of atoms with the Pearson correlation coefficient in Fig. 4. It is not surprising to see a non-negligible correlation between properties and molecular sizes, since the sizes of molecules typically constrain their possibility over the chemical space. To ensure a fair comparison, we stick to the molecular space with similar size to the reference. For further comparison among different model variants, we report the molecular properties under different sizes in Table 5. Results show that our method consistently achieves the highest success rate, demonstrating its robust optimization ability even in an Out-of-Distribution (OOD) scenario.
@@ Appendix F Full Optimization Results
  • Pocket2Mol (Peng et al., 2022) generates one atom with bond at a time via an E(3)-equivariant network. It predicts frontier atoms to expand, alleviating the efficiency problem in sampling.
  • FLAG (Zhang et al., 2023) is a fragment-based model that assembles the generated fragments via predicted coordinates and torsion angles.
@@
  • IPDiff (Huang et al., 2024) pretrains an affinity predictor, and utilizes this predictor to extract features that augments the conditioning of diffusion generative process.
  • MolCRAFT (Qu et al., 2024) employs Bayesian Flow Network for molecular design with an advanced sampling strategy, showing notable improvement upon diffusion counterparts.
  • AutoGrow4 (Spiegel & Durrant, 2020) is an evolutionary algorithm that uses genetic algorithm to optimize 1D SMILES, with additional docking simulation. Starting from the initial seed molecule, AutoGrow4 iteratively conducts mutations and crossovers, then makes oracle calls for docking feedback, and keeps the top-scoring molecules in the end.
@@
  • DecompOpt (Zhou et al., 2024a) trains a conditional generative model on decomposed fragments besides binding pocket, following the style of DecompDiff. The optimization is done by iteratively re-sampling in the 3D diffusion latent space given the top $K$ arms ranked by oracle functions as updated fragment condition input.
  • TAGMol (Dorna et al., 2024) exerts gradient-based property guidance over the pretrained TargetDiff backbone, and the gradient is enabled only in the continuous diffusion process for coordinates.
@@
Besides the common evaluation metrics such as binding affinities calculated by Autodock Vina (Eberhardt et al., 2021) and QED, SA by RDKit, we elaborate other metrics as follows:
  • Diversity measures the diversity of generated molecules for each binding site. Following SBDD convention (Luo et al., 2021), it is based on Tanimoto similarity over Morgan fingerprints, and averaged on 100 test proteins.
  • Connected Ratio is the ratio of complete molecules among all, i.e. with only one connected component.
  • Lipinski enumerates Lipinski rule of five (Lipinski et al., 1997) and checks how many rules are satisfied. These rules are typically seen as empirical reference that helps to predict whether the molecule is likely to be orally bioavailable.
@@
  • Steric Clash calculates the number of clashes between generated ligand and protein surface, where clashing means the distance of ligand and protein atoms are within a certain threshold. This reveals the stability of complex to some extend, yet it does not strictly mean violation of physical constraint since the protein is not that rigid and might also go through spatial rearrangement upon binding, as noted by Harris et al. (2023).
  • Redocking RMSD reports the percentage of molecules with an RMSD between generated and Vina redocked poses lie within the range of 2Å, which suggests the binding mode remains consistent after redocking.

### v4

# Empower Structure-Based Molecule Optimization with Gradient Guided Bayesian Flow Networks

@@ Abstract
Structure-based molecule optimization (SBMO) aims to optimize molecules with both continuous coordinates and discrete types against protein targets. A promising direction is to exert gradient guidance on generative models given their remarkable success in images, but it is challenging to guide discrete data and risks inconsistencies between modalities. To this end, we leverage a continuous and differentiable space derived through Bayesian inference, presenting Molecule Joint Optimization (MolJO), the gradient-based SBMO framework that facilitates joint guidance signals across different modalities while preserving SE(3)-equivariance. We introduce a novel backward correction strategy that optimizes within a sliding window of the past, allowing for a seamless trade-off between exploration and exploitation during optimization. MolJO achieves state-of-the-art performance on CrossDocked2020 benchmark (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), more than 4×4$\times$4 × improvement in Success Rate compared to the gradient-based counterpart, and 2×2$\times$2 × “Me-Better” Ratio as high as 3D baselines. Furthermore, we extend MolJO to a wide range of settings, including multi-objective optimization and challenging tasks in drug design such as R-group optimization and scaffold hopping, further underscoring its versatility. Code is available at https://github.com/AlgoMole/MolCRAFT.
@@ 1 Introduction
Structure-based drug design (SBDD) plays a critical role in drug discovery by identifying three-dimensional (3D) molecules that are favorable for protein targets (Isert et al., 2023). While recent SBDD focuses on the initial identification of potential drug candidates, these compounds must undergo a series of further modifications for optimized properties, a process that is both complex and time-consuming (Hughes et al., 2011). Therefore, structure-based molecule optimization (SBMO) has garnered increasing interest in real-world drug design (Zhou et al., 2024a), emphasizing the practical need to optimize for specific therapeutic criteria.
@@
In this paper, we address the multi-modality challenge for gradient guidance by leveraging a continuous and differential space, representing an aggregation of noisy samples from the data space derived through Bayesian inference (Graves et al., 2023). We design MolJO (Molecule Joint Optimization), a principled, end-to-end differentiable framework that enables gradient-based optimization of continuous and discrete variables. We introduce a novel sampling strategy called backward correction, enhancing the alignment of gradients over different steps. By maintaining a sliding window of past history for optimization, the backward correction strategy enforces explicit dependency on the past, effectively alleviating the issue of inconsistencies. Moreover, it balances the exploration of molecular space with the exploitation of better-aligned guidance signals, offering a flexible trade-off.
@@ 6 Conclusion
We present MolJO, the joint gradient-based SE(3)-equivariant framework within Bayesian Flow Networks to solve the structure-based molecule optimization problems, which only requires differentiable energy functions instead of expensive oracle simulations. The general framework further equips gradient-based optimization method with backward correction strategy, offering a flexible trade-off between exploration and exploitation. Experiments show that MolJO is able to improve the binding affinity of molecules by establishing more key interactions and enhance drug-likeness and synthesizability, achieving state-of-the-art performance on CrossDocked2020 (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), together with 4×4$\times$4 × improvement compared to gradient-based counterpart and 2×2$\times$2 × “Me-Better” Ratio as other 3D baselines.
@@ Acknowledgments
This work is supported by the Natural Science Foundation of China (Grant No. 62376133) and sponsored by Beijing Nova Program (20240484682) and the Wuxi Research Institute of Applied Technologies, Tsinghua University (20242001120).
@@ Appendix A Overview of Bayesian Flow Networks
In this section, we provide further explanation of Bayesian Flow Networks (BFNs) that are designed to model the generation of data through a process of message exchange between a “sender” and a “receiver” (Graves et al., 2023). The fundamental elements include the sender distribution for the sender, and input distribution, output distribution, receiver distribution for the receiver.
@@
This process is framed within the context of Bayesian inference, where the sender distribution is a factorized distribution $p_{S}(\mathbf{y}|\mathbf{m},\alpha_{t})$ that introduces noise to each dimension of the data $\mathbf{m}$ and sends it to the receiver. The receiver observes $\mathbf{y}$, has access to the noisy channel with accuracy $\alpha$ at timestep $t$, and compares it with its own receiver distribution $p_{R}(\mathbf{y}|\bm{\theta},\mathbf{p};t)$ based on its current belief of the parameters $\bm{\theta}$, the timestep and any conditional input such as protein pocket $\mathbf{p}$.
@@
The generative process for the receiver begins with a prior distribution (referred to as input distribution $p_{I}(\mathbf{m}|\bm{\theta})=\prod_{d=1}^{N}p_{I}(m^{(d)}|\theta^{(d)})$ that is also factorized for $N$-dimensional data) that defines its initial belief about the data. For continuous data, the prior can be chosen as a Gaussian (Song et al., 2024; Qu et al., 2024), or another distribution such as von Mises distribution (Wu et al., 2025), while for discrete data, the prior is modeled as a uniform categorical distribution. Then, the receiver uses its belief $\bm{\theta}$ with the help of a neural network to model the inter-dependency among dimensions and compute the output distribution $p_{O}(\hat{\mathbb{m}}|\bm{\theta},\mathbf{p};t)$, which represents its estimate of the possible reconstructions of the original data $\mathbf{m}$, and is used to construct the receiver distribution (Eq. 2).
@@
The Bayesian update function in Eq. 1 defines how the prior belief $\bm{\theta}_{0}$ is updated to the conjugate posterior $\bm{\theta}_{t}$. Ideally, the update requires aggregating all possible noisy latents $\mathbf{y}$ from the sender distribution $p_{S}$. However, during the actual generative process, only the receiver distribution $p_{R}$ is available, which leaves an exposure bias, and different approximations determine different forms of mapping $\theta=f(\mathbf{y})$ to the posterior, showcasing the flexibility in the design space of BFN.
@@
Through the iterative communication between sender and receiver, the receiver progressively updates its belief of the underlying parameters, and training is achieved by minimizing the divergence between the sender and receiver distributions (Eq. 3). This is analogous to the way Bayesian inference works in parameter estimation: as more noisy data $\mathbf{y}$ is observed, the receiver’s posterior belief about the data $\mathbf{m}$ becomes increasingly accurate, which implies the reconstruction would be made easier. It is shown that BFN for Gaussian priors can be sampled from the view of SDE (Xue et al., 2024), but this type of sampling displays distinct behaviors empirically, possibly due to discretization errors.
@@ Appendix B Sender Distribution for Discrete Data
While true discrete data can be viewed as a sharp one-hot distribution, it can be further relaxed by a factor $\omega\in[0,1]$ into a Categorical distribution defined by the probability $p(k^{(d)}|\mathbf{v}^{(d)};\omega)=\frac{1-\omega}{K}+\omega\delta_{k^{(d)}\mathbf{v}^{(d)}}$ for $k$ from $1$ to $K$ along the $d$-th dimension, where $\delta$ is the Kronecker delta function.
Instead of focusing on the density or sampling from it once, note that the counts $c$ of observing each class in $m$ independent draws follow a multinomial distribution, namely $c\sim\text{Multi}(m,p)$. Dropping the superscripts, Graves et al. (2023) derives the following conclusions:
@@
Thus, it naturally follows that such noisy $\mathbf{y}^{v}\sim\mathcal{N}(\mathbf{y}^{v}|\alpha(K\mathbf{e}_{\mathbf{v}}-\mathbb{1}),\alpha K\mathbb{I})$.
@@ C.1 Proof of Guided Bayesian Update Distribution
where some parentheses and protein pocket condition $\mathbf{p}$ have been omitted for brevity.
@@
Eq. 19, together with the lemma above, leads to Proposition 4.1.
@@ C.2 Proof of Equivariance
For any orthogonal matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ such that $\mathbf{R}^{\top}\mathbf{R}=\mathbb{I}$, it is easy to see that the prior $\bm{\theta}^{x}_{0}=\mathbb{0}$ is O(3)-invariant. Given that $\hat{\mathbf{x}}\sim p_{O}(\hat{\mathbf{x}}~{}|~{}\bm{\Phi}(\bm{\theta},\mathbf{p},t))$ and the equivariance of $\bm{\Phi}$, it suffices to prove the invariant likelihood for the transition kernel.
Given the parameterization of the pretrained energy function $E(\bm{\theta},\mathbf{p},t)$ is SE(3)-equivariant, then the gradient $\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta})=-\nabla_{\bm{\theta}^{x}}E(\bm{\theta},\mathbf{p},t_{i})$ is also equivariant according to Bao et al. (2022).
@@
It also applies to cases where $i>k$, as we can recurrently view the starting point of backward corrected history $\bm{\theta}^{x}_{i-k}$ as the new O(3)-invariant prior $\bm{\theta}^{x}_{0}$ and iteratively make the above derivation. ∎
@@ D.1 Model Details
To enable a differentiable oracle function, we additionally train the energy function based on the molecules and their properties (Vina Score, QED, SA) in CrossDocked dataset (Francoeur et al., 2020) by minimizing the squared loss for property $c$ over the data distribution $p_{\mathrm{data}}$:
@@
The input parameters to all energy functions belong to the parameter space defined by $\beta_{1}=1.5$ for atom types, $\sigma_{1}=0.03$ for atom coordinates, $n=1000$ discrete steps. The energy network is parameterized with the same model architecture as TargetDiff (Guan et al., 2022), i.e. kNN graphs with $k=32$, $N=9$ layers with $d=128$ hidden dimension, 16-headed attention, and the same featurization, i.e. protein atoms (H, C, N, O, S, Se) and ligand atoms (C, N, O, F, P, S, Cl). For training, the Adam optimizer is adopted with learning rate 0.0005, batch size is set to 8. The training takes less than 8 hours on a single RTX 3090 and converges within 5 epochs.
@@
To sample via guided Bayesian flow, we set the sample steps to 200, and the guidance scale to 50. For the combination of different objectives, we simply take an average of different gradients.
@@ D.2 Task Details
Lead optimization cases involve retaining the scaffold while redesigning the remaining R-groups, usually when the scaffold forms desirable interactions with the protein and anchors the binding mode, and the remaining parts need further modifications to secure this pattern and enhance binding affinity. Following Polykovskiy et al. (2020), we employ RDKit for fragmentation and atom annotation with R-group or Bemis-Murcko scaffold.
@@
Unlike R-group design, scaffold hopping involves redesigning the scaffold for a given molecule while keeping its core functional groups, for example, to overcome the patent protection for a known drug molecule while retaining pharmaceutical activity. This is a technically more challenging task for generative models, because the missing parts they need to fulfill are generally larger than those in R-group design, and the hopping is usually subject to more chemical constraints. We construct scaffold hopping as a dual problem to R-group optimization using Bemis-Murcko scaffolding annotation, although it does not need to be so.
@@ Appendix E Effect of Molecular Size on Properties
The size of molecules is found to have a notable impact on molecular properties, including Vina affinities (Qu et al., 2024). We quantify the relationship and plot the distribution of molecular properties w.r.t. the number of atoms with the Pearson correlation coefficient in Fig. 4. It is not surprising to see a non-negligible correlation between properties and molecular sizes, since the sizes of molecules typically constrain their accessible chemical space. To ensure a fair comparison, we adhere to the molecular space with similar size to the reference. For further comparison among different model variants, we report the molecular properties under different sizes in Table 5. Results show that our method consistently achieves the highest success rate, demonstrating its robust optimization ability even in an Out-of-Distribution (OOD) scenario.
@@ Appendix F Full Optimization Results
  • Pocket2Mol (Peng et al., 2022) generates one atom and its bond at a time via an E(3)-equivariant network. It predicts frontier atoms to expand, alleviating the efficiency problem in sampling.
  • FLAG (Zhang et al., 2023) is a fragment-based model that assembles the generated fragments using predicted coordinates and torsion angles.
@@
  • IPDiff (Huang et al., 2024) pretrains an affinity predictor, and utilizes this predictor to extract features that augment the conditioning of the diffusion generative process.
  • MolCRAFT (Qu et al., 2024) employs Bayesian Flow Networks for molecular design with an advanced sampling strategy, showing notable improvement upon diffusion counterparts.
  • AutoGrow4 (Spiegel & Durrant, 2020) is an evolutionary algorithm that uses a genetic algorithm to optimize 1D SMILES with docking simulation. Starting from the initial seed molecule, AutoGrow4 iteratively conducts mutations and crossovers, then makes oracle calls for docking feedback, and retains the top-scoring molecules in the end.
@@
  • DecompOpt (Zhou et al., 2024a) trains a conditional generative model on decomposed fragments and the binding pocket, following the style of DecompDiff. The optimization is done by iteratively resampling in the 3D diffusion latent space given the top $K$ arms ranked by oracle functions as updated fragment condition input.
  • TAGMol (Dorna et al., 2024) exerts gradient-based property guidance on the pretrained TargetDiff backbone, and the gradient is enabled only in the continuous diffusion process for coordinates.
@@
Besides the common evaluation metrics such as binding affinities calculated by Autodock Vina (Eberhardt et al., 2021) and QED, SA calculated by RDKit, we elaborate other metrics as follows:
  • Diversity measures the diversity of generated molecules for each binding site. Following SBDD convention (Luo et al., 2021), it is based on Tanimoto similarity over Morgan fingerprints, and averaged across 100 test proteins.
  • Connected Ratio is the ratio of complete molecules overall, i.e. with only one connected component.
  • Lipinski enumerates the Lipinski rule of five (Lipinski et al., 1997) and checks how many are satisfied. These rules are typically seen as an empirical reference that helps to predict whether the molecule is likely to be orally bioavailable.
@@
  • Steric Clash calculates the number of clashes between generated ligand and protein surface, where clashing means the distance of ligand and protein atoms is within a certain threshold. This reveals the stability of complex to some extend, yet it does not strictly mean a violation of physical constraints, since the protein is not overly rigid and might also undergo spatial rearrangement upon binding, as noted by Harris et al. (2023).
  • Redocking RMSD reports the percentage of molecules with an RMSD between generated and Vina redocked poses lying within the range of 2Å, which suggests the binding mode remains consistent after redocking.